Find the value of sine 15 degree × sine 75 degree ?

Grade:11

1 Answers

yathartha gupta
6 years ago
= 2/2sin15×sin75= 1/2[cos(15+75)-cos(75-15)]= 1/2[cos90-cos60]= 1/2[0-1/2]= -1/4~~~~~~~~~~~~~~~~~~~ANSWER